OpenStudy (anonymous):

Suppose you are given that tan(theta)=3. The hypotenuse=9, so find area of triangle. Please Help!

OpenStudy (anonymous):

\[a ^{2}+3a ^{2}=81\]\[b=3a\]

OpenStudy (anonymous):

that will find your sides a,b and then you can get the area.

OpenStudy (anonymous):

\[\tan \theta = \frac{ opposite }{ adjacent } = 3\]if we let opposite be b and adjacent be a we have \[\frac{ b }{ a }=3=\frac{ 3a }{ a }\]

OpenStudy (anonymous):

That is how I arrived at my equation above. now solve for a, then you can get b, therefore the area

OpenStudy (anonymous):

Are you good from here?

OpenStudy (anonymous):